OpenStudy (anonymous):

I just want to make sure I'm doing this right: find dy/dx by implicit differentiation x^3-xy+y^2=4

OpenStudy (anonymous):

\[3x ^{2}-\left(\begin{matrix}dy \\dx\end{matrix}\right)+2y \left(\begin{matrix}dy \\dx\end{matrix}\right)=0\]

OpenStudy (anonymous):

right?

OpenStudy (ranga):

For the middle term apply product rule. The last two terms are correct.

OpenStudy (kenljw):

x^3-xy+y^2=4 3x^2 - y -x(dy/dx) + 2y(dydx) = 0 dy/dx = (3x^2 - y)/(x - 2y)

OpenStudy (anonymous):

Oh! ok thank you!